Find all distinct eigenvalues of A. Then find the basic eigenvectors of A corresponding to each eigenvalue. For each eigenvalue, specify the number of basic eigenvectors corresponding to that eigenvalue, then enter the eigenvalue followed by the basic eigenvectors corresponding to that eigenvalue 2 12 6 A 0 -14 -8 0 24 14 Number of distinct eigenvalues: 1 Number of Vectors: 1 030
Answer #1

Solution: The characteristic equation is

\small \lambda ^{3}-\left ( \mathrm{sum\: of\: diagonal\: elements }\right )\lambda ^{2}+\left ( \mathrm{sum\: of\: minor\: of\: diagonals }\right )\lambda -\left | A \right |=0

\small \therefore \lambda ^{3}-2\lambda ^{2}-4\lambda +8=0

\small \therefore \left ( \lambda -2 \right )\left ( \lambda ^{2}-4 \right )=0

\small \therefore \lambda =2\: ,\: 2\: ,\: -2

For   \small \lambda _{1}=2\: \: \: \: \: \: \left [ A-\lambda _{1}I \right ]X=0

\small \begin{bmatrix} 0 &12 &6 \\ 0 &-16 &-8 \\ 0& 24 & 12 \end{bmatrix}\begin{bmatrix} x\\ y\\ z\\ \end{bmatrix}=\begin{bmatrix} 0\\ 0\\ 0\\ \end{bmatrix}

By \small R_{1}\times \frac{1}{6}\: \: ,\: \: R_{2}\times \frac{-1}{8}\: \: ,\: \: R_{3}\times \frac{1}{12}

\small \small \begin{bmatrix} 0 &2 &1 \\ 0 &2 &1 \\ 0& 2 & 1 \end{bmatrix}\begin{bmatrix} x\\ y\\ z\\ \end{bmatrix}=\begin{bmatrix} 0\\ 0\\ 0\\ \end{bmatrix}

By \small R_{2}-R_{1}\: \: ,\: \: R_{3}-R_{1}

0 210 0001 11,1

The rank of \small A   is \small 1< 3   the number of unknown .

\small \therefore The system has nontrivial solution. The number of linearly independent solution is \small 3-1=2

\small \therefore 0\cdot x+2y+z=0

Put \small x=s\: ,\: \: y=t

\small \therefore z=-2t

\small \therefore The solution space is

\small X=\begin{bmatrix} s\\ t\\ -2t\\ \end{bmatrix}=s\begin{bmatrix} 1\\ 0\\ 0\\ \end{bmatrix}+t\begin{bmatrix} 0\\ 1\\ -2\\ \end{bmatrix}

\small \therefore X_{1}=\begin{bmatrix} 1\\ 0\\ 0\\ \end{bmatrix}

\small \therefore X_{2}=\begin{bmatrix} 0\\ 1\\ -2\\ \end{bmatrix}

For   \small \lambda _{2}=-2\: \: \: \: \: \: \left [ A-\lambda _{2}I \right ]X=0

\small \begin{bmatrix} 4 &12 &6 \\ 0 &-12 &-8 \\ 0& 24 & 16 \end{bmatrix}\begin{bmatrix} x\\ y\\ z\\ \end{bmatrix}=\begin{bmatrix} 0\\ 0\\ 0\\ \end{bmatrix}

By \small R_{1}\times \frac{1}{2}\: \: ,\: \: R_{2}\times \frac{-1}{4}\: \: ,\: \: R_{3}\times \frac{1}{8}

\small \begin{bmatrix} 2 &6 &3 \\ 0 &3 &2 \\ 0& 3 & 2 \end{bmatrix}\begin{bmatrix} x\\ y\\ z\\ \end{bmatrix}=\begin{bmatrix} 0\\ 0\\ 0\\ \end{bmatrix}

By \small R_{3}-R_{2}

\small \begin{bmatrix} 2 &6 &3 \\ 0 &3 &2 \\ 0& 0 & 0 \end{bmatrix}\begin{bmatrix} x\\ y\\ z\\ \end{bmatrix}=\begin{bmatrix} 0\\ 0\\ 0\\ \end{bmatrix}

The rank of \small A   is \small 2< 3   the number of unknown .

\small \therefore The system has nontrivial solution. The number of linearly independent solution is \small 3-2=1

\small 3y+2z=0

\small 2x+6y+3z=0

Put   \small z=3t

\small \therefore y=-2t   and   \small x=\frac{3}{2}t

\small \therefore The solution space is

\small X_{3}=\begin{bmatrix} \frac{3}{2}\\ -2t\\ 3t\\ \end{bmatrix}=t\begin{bmatrix} \frac{3}{2}\\ -2\\ 3\\ \end{bmatrix}

Number of distinct eigenvalues: \small {\color{Red} 2}

Number of vectors: \small {\color{Magenta} 3}

\small {\color{DarkGreen} 2:\left \{ \begin{bmatrix} 1\\ 0\\ 0\\ \end{bmatrix}\: \: ,\: \: \begin{bmatrix} 0\\ 1\\ -2\\ \end{bmatrix} \right \}}

\small {\color{DarkRed} -2:\left \{ \begin{bmatrix} \frac{3}{2}\\ -2\\ 3\\ \end{bmatrix} \right \}}